### v2

# First Place Solution to the ECCV 2024 BRAVO Challenge: Evaluating Robustness of Vision Foundation Models for Semantic Segmentation

Abstract
In this report, we present the first place solution to the ECCV 2024 BRAVO Challenge, where a model is trained on Cityscapes and its robustness is evaluated on several out-of-distribution datasets. Our solution leverages the powerful representations learned by vision foundation models, by attaching a simple segmentation decoder to DINOv2 and fine-tuning the entire model. This approach outperforms more complex existing approaches, and achieves first place in the challenge. Our code is publicly available [^1] https://github.com/tue-mps/benchmark-vfm-ss.
1 Introduction
The ECCV 2024 BRAVO Challenge “aims to benchmark semantic segmentation models on urban scenes undergoing diverse forms of natural degradation and realistic-looking synthetic corruption" [17]. In this report, we present our solution for Track 1 of this challenge, where a model is trained on a single labeled semantic segmentation dataset, Cityscapes [4], and evaluated on a range of other datasets with out-of-distribution image conditions and semantic content, to evaluate the robustness of the model.
When trained on a single dataset like Cityscapes, vanilla semantic segmentation models typically only perform well on data that is similar to this training dataset. Under a distribution shift – e.g., changing weather conditions, different geography or image distortion – the performance drops. To address this problem, we do not propose a new algorithmic approach or model design, but instead aim to leverage the power of Vision Foundation Models (VFMs). VFMs are pre-trained on broad datasets, allowing them to learn powerful representations for a large variety of downstream tasks [12, 8, 14, 21, 1, 18]. As these VFMs learn representations from such diverse data, captured under a wide array of conditions, we hypothesize that they are perfectly suited for robust semantic segmentation with simple, vanilla fine-tuning.
In this report, we present our simple meta-approach, and evaluate several configurations. In our default configuration, we use a DINOv2 [12] VFM, attach a simple linear decoder for segmentation, and fine-tune the entire model. We assess the impact of using different model sizes, patch sizes, pre-training strategies and segmentation decoders. With our simple approach, we significantly outperform existing specialist models and obtain the first place in the challenge. Moreover, we make some new observations which could be of interest to future work.
Figure 1: Our meta-approach. We take a pre-trained Vision Foundation Model (VFM), attach a simple segmentation decoder, and fine-tune the entire model for semantic segmentation. The segmentation decoder outputs both the per-pixel classification predictions and the associated confidence scores.
2 Method
The concept of our approach is to fine-tune pre-trained Vision Foundation Models (VFMs) for semantic segmentation, leveraging the robust representations these models provide. The meta-approach is visualized in Figure 1. Given a pre-trained VFM, we attach an off-the-shelf segmentation decoder, and fine-tune the entire model for semantic segmentation. We evaluate this meta-architecture in several different configurations.
Our primary solution, which achieves first place in the challenge, uses the DINOv2 VFM [12]. We selected DINOv2 due to its demonstrated effectiveness in domain-generalized semantic segmentation for urban scenes [10, 7]. DINOv2, built upon the Vision Transformer (ViT) architecture [6], is pre-trained using self-supervised learning on a vast, curated dataset. We conduct experiments using all available sizes of DINOv2.
For our default segmentation decoder, we use a simple linear layer that transforms the patch-level features $\mathbf{F}\in\mathbb{R}^{{E}\times\frac{H}{P}\times\frac{W}{P}}$ into segmentation logits $\mathbf{L}\in\mathbb{R}^{{C}\times\frac{H}{P}\times\frac{W}{P}}$, where $H$ and $W$ represent the height and width of the input image, $P$ denotes the patch size, $E$ is the feature dimensionality, and $C$ is the number of classes in the dataset. We opt for a linear layer because we hypothesize that a more advanced decoder would provide minimal additional benefit given the already strong representations learned by the VFM. Additionally, a more advanced decoder could increase overfitting to the training distribution.
To assess the impact of different design choices of our default configuration, we compare it with alternative configurations. First, to evaluate the impact of large-scale pre-training with DINOv2, we make a configuration with a DeiT-III [16] ViT pre-trained on ImageNet-1K [5] and fine-tune it on Cityscapes. Second, we train a model with a Mask2Former decoder [3] to assess the impact of a more advanced decoder. Finally, where the default patch size used is $16\times 16$, we also experiment with the more computationally expensive $8\times 8$ patch size.
2.1 Training
When training the model with a linear decoder, we bilinearly upsample the segmentation logits $\mathbf{L}\in\mathbb{R}^{{C}\times\frac{H}{P}\times\frac{W}{P}}$ to $\mathbf{L}^{\prime}\in\mathbb{R}^{{C}\times{H}\times{W}}$, and then apply a categorical cross-entropy loss to these logits and the semantic segmentation ground truth to fine-tune the model.
When using Mask2Former, the decoder outputs a set of mask logits $\mathbf{M}\in\mathbb{R}^{{N}\times{\frac{H}{P}}\times{\frac{W}{P}}}$ and corresponding class logits $\mathbf{C}\in\mathbb{R}^{{N}\times{({C}+1})}$, where $N$ is the number of masks and $\mathbf{C}$ includes an additional “no-object” class. During training, following Mask2Former, these mask and class logits are matched to the ground truth using bipartite matching. The predicted masks are then supervised with a cross-entropy loss and a Dice loss, and the predicted classes are supervised with a categorical cross-entropy loss.
2.2 Testing
During inference with the linear decoder, we compute per-pixel class confidence scores by applying a softmax function to the upsampled class logits $\mathbf{L}^{\prime}$. For each pixel, the predicted class is the one with the highest confidence score, and we also output this confidence score.
During inference with the Mask2Former decoder [3], we first bilinearly upsample the mask logits $\mathbf{M}\in\mathbb{R}^{{N}\times{\frac{H}{P}}\times{\frac{W}{P}}}$ to the original resolution, resulting in $\mathbf{M}^{\prime}\in\mathbb{R}^{{N}\times{H}\times{W}}$. We then apply a sigmoid function to $\mathbf{M}^{\prime}$ to obtain the mask scores: $\mathbf{P}_{\textrm{M}}=\texttt{sigmoid}(\mathbf{M}^{\prime})$.
The class logits $\mathbf{C}\in\mathbb{R}^{{N}\times{({C}+1})}$ are converted to class scores using a softmax function, excluding the “no object" class: $\mathbf{P}_{\textrm{C}}=\texttt{softmax}(\mathbf{C})[...,:-1]$.
The overall per-pixel class confidence scores $\mathbf{P}^{\prime}\in\mathbb{R}^{C\times H\times W}$ are computed by multiplying the mask scores with the class scores across all masks. Specifically, for each class $c$ and pixel $(h,w)$, we have:
$\mathbf{P}^{\prime}_{c,h,w}=\sum_{n=1}^{N}\mathbf{P}_{\textrm{M}_{n,h,w}}\cdot\mathbf{P}_{\textrm{C}_{n,c}}$ (1)
For each pixel, the predicted class is the one with the highest value in $\mathbf{P}^{\prime}$, and we also output this maximum value as the confidence score.
3 Experimental Setup
3.1 Datasets
Track 1 of the 2024 BRAVO Challenge focuses on single-domain training, assessing the robustness of models trained with limited supervision and geographical diversity against real-world distribution shifts. Only the Cityscapes dataset [4] is permitted for training. This dataset features homogeneous urban scenes with fine-grained annotations for 19 semantic classes.
The BRAVO benchmark dataset, used for evaluation, consists of six subsets:
  • ACDC [15]: Real scenes captured in adverse weather conditions, such as fog, rain, night, and snow.
  • SMIYC [2]: Real scenes featuring out-of-distribution (OOD) objects that are rarely encountered on the road.
  • Out-of-context [9]: Augmented scenes with random backgrounds, generated by replacing the backgrounds of 329 validation images from Cityscapes with random ones.
  • Synflare [20]: Augmented scenes with synthesized light flares, produced by adding random light flares to 308 validation images from Cityscapes.
  • Synobjs [11]: Augmented scenes with inpainted synthetic OOD objects, created by adding 26 different OOD objects to 656 validation images from Cityscapes.
  • Synrain [13]: Augmented scenes with synthesized raindrops on the camera lens, generated by augmenting 500 validation images from Cityscapes.
3.2 Evaluation metrics
The ECCV 2024 BRAVO Challenge evaluates methods based on a variety of metrics to assess their performance in semantic segmentation and out-of-distribution (OOD) detection. The metrics are grouped into three categories: semantic metrics, OOD metrics, and summary metrics. The semantic metrics evaluate the overall quality of the semantic segmentation predictions, while the OOD metrics assess the model’s ability to detect OOD objects. The summary metrics combine the semantic and OOD metrics to provide an overall ranking of the methods.
Semantic metrics.
The semantic metrics are computed on all subsets, except SMIYC, for valid pixels only. Valid pixels are those not “invalidated” by extreme uncertainty, such as pixels obscured by the brightest areas of a flare or covered by an OOD object. The semantic metrics include:
  • Mean Intersection over Union (mIoU): Measures the proportion of correctly labeled pixels among all pixels. It is the only semantic metric that does not rely on prediction confidence. Higher mIoU values indicate better segmentation accuracy.
  • Expected Calibration Error (ECE): Quantifies the difference between predicted confidence and actual accuracy. Lower ECE values indicate better calibration of the model’s confidence scores.
  • Area Under the ROC Curve (AUROC): Represents the area under the curve plotting the true positive rate against the false positive rate, using the predicted confidence level to rank pixels. Higher AUROC values indicate better ability to distinguish between correct and incorrect predictions.
  • False Positive Rate at 95% True Positive Rate (FPR@95): Measures the false positive rate when the true positive rate is 95% computed in the ROC curve above. Lower FPR@95 values indicate better robustness against false positives.
  • Area Under the Precision-Recall Curve (AUPR): Represents the area under the curve plotting precision against recall, using the predicted confidence level to rank pixels. Higher AUPR-Success and AUPR-Error values indicate better ability to identify correct and incorrect predictions, respectively.
OOD metrics.
The OOD metrics are computed on the SMIYC and Synobjs subsets only. Invalid pixels in these datasets are those that are obscured by OOD objects. The OOD metrics quantify whether the model attributes, as expected, less confidence to the invalid pixels. The OOD metrics include:
  • Area Under the Precision-Recall Curve (AUPRC): AUPRC, over the binary criterion of a pixel being invalid, ranked by the reversed predicted confidence level for the pixel.
  • Area Under the ROC Curve (AUROC): Area Under the ROC Curve, over the binary criterion of a pixel being invalid, ranked by the reversed predicted confidence level for the pixel.
  • False Positive Rate at 95% True Positive Rate (FPR@95): False Positive Rate when True Positive Rate is 95% computed in the ROC curve above.
Summary metrics.
The summary metrics include:
  • Semantic: The harmonic mean of all semantic metrics, where ECE and FPR@95 are reversed.
  • OOD: The harmonic mean of all OOD metrics, where FPR@95 is reversed.
  • BRAVO: The harmonic mean of the semantic and OOD harmonic means, serving as the official ranking metric for the challenge.
3.3 Implementation Details
We use the following models from the timm library [19] to initialize the VFM:
  • deit3_small_patch16_224.fb_in1k;
  • vit_small_patch14_dinov2;
  • vit_base_patch14_dinov2;
  • vit_large_patch14_dinov2;
  • vit_giant_patch14_dinov2.
The models are fine-tuned for 40 epochs using two A6000 GPUs, with a batch size of 1 per GPU and gradient accumulation over 8 steps, resulting in an effective batch size of 16. Our implementation follows the details provided in [10]. Notably, the learning rate for the VFM weights is set to be 10$\times$ smaller than the overall learning rate, as this configuration empirically yields better results. For the Mask2Former decoder, we employ a variant specifically adapted for use with a single-scale ViT encoder, as introduced in [10].
4 Results
Method BRAVO $\uparrow$ Semantic $\uparrow$ OOD $\uparrow$
DINOv2, ViT-L, 8x8 patch size, linear decoder 77.9 69.8 88.1
DINOv2, ViT-L, 16x16 patch size, linear decoder 77.2 70.8 84.8
DINOv2, ViT-g, 16x16 patch size, linear decoder 76.1 70.0 83.4
DINOv2, ViT-B, 16x16 patch size, linear decoder 75.5 70.5 81.4
DINOv2, ViT-S, 16x16 patch size, linear decoder 69.9 69.1 70.6
PixOOD YOLO (="Model Selection") 67.8 57.1 83.5
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 64.5 49.7 92.1
Model selection 63.5 69.4 58.5
PixOOD w/ ResNet-101 DeepLab 61.2 58.7 64.0
Ensemble C 61.1 64.3 58.2
Ensemble A 59.9 67.3 53.9
PixOOD w/ DeepLab Decoder 59.4 46.1 83.5
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 54.1 62.8 47.6
PixOOD 53.5 40.4 79.1
SegFormer-B5 47.1 45.3 49.2
ObsNet-R101-DLv3plus 45.3 51.5 40.5
Mask2Former-SwinB 37.7 27.7 59.2
Physically Feasible Semantic Segmentation 33.6 66.3 22.5
Table 1: BRAVO index. The official ranking metric, calculated as the harmonic mean of the semantic and OOD harmonic means for each method. Grayed-out methods are submitted by other participants.
BRAVO index. The official ranking metric, calculated as the harmonic mean of the semantic and OOD harmonic means for each method, is shown in Table 1. Our best-performing model, DINOv2 with a ViT-L/8 backbone and a linear decoder, achieves the highest BRAVO index of 77.9, which is +10.1 higher than the next best method not submitted by us, which uses multiple different specialist models, i.e., PixOOD YOLO (="Model Selection").
Method ACDC $\uparrow$ SMIYC $\uparrow$ Out-of-context $\uparrow$ Synflare $\uparrow$ Synobjs $\uparrow$ Synrain $\uparrow$
DINOv2, ViT-L, 8x8 patch size, linear decoder 67.3 89.9 71.0 72.7 76.7 73.9
DINOv2, ViT-L, 16x16 patch size, linear decoder 69.4 89.3 70.4 72.4 75.1 73.8
DINOv2, ViT-g, 16x16 patch size, linear decoder 67.7 88.2 71.0 73.2 74.7 73.2
DINOv2, ViT-B, 16x16 patch size, linear decoder 68.5 87.9 71.2 72.8 74.0 73.0
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.9 83.1 70.2 70.6 68.6 72.9
PixOOD YOLO (="Model Selection") 55.7 74.9 64.9 65.0 56.4 63.3
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 49.1 94.4 40.9 53.9 64.3 60.1
Model selection 66.6 70.7 73.4 77.7 58.9 76.4
PixOOD w/ ResNet-101 DeepLab 55.7 53.2 64.9 65.0 63.2 63.3
Ensemble C 64.8 63.1 61.7 69.6 58.1 64.6
Ensemble A 66.4 58.7 66.0 74.0 58.9 69.4
PixOOD w/ DeepLab Decoder 48.4 74.9 56.2 43.6 56.4 36.4
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 58.8 50.1 65.1 71.4 58.5 65.6
PixOOD 40.8 66.7 50.0 48.4 49.4 36.4
SegFormer-B5 45.5 60.5 39.1 50.6 41.7 51.5
ObsNet-R101-DLv3plus 50.8 38.7 50.7 50.9 50.2 52.0
Mask2Former-SwinB 28.6 64.5 23.5 41.1 29.3 26.7
Physically Feasible Semantic Segmentation 63.6 20.9 68.5 68.1 43.1 71.0
Table 2: Subset harmonic means. Harmonic means of semantic and OOD metrics for each subset in the BRAVO benchmark dataset, computed for each method. Grayed-out methods are submitted by other participants.
Subset harmonic means. The harmonic means of semantic and OOD metrics for each subset in the BRAVO benchmark dataset are shown in Table 2. Although our DINOv2-based models perform well across most subsets, they do not consistently outperform all other methods. Specifically, on the Out-of-context, Synobjs, and Synrain subsets, the method named Model selection outperforms our models. Interestingly, compared to other methods, our DINOv2-based models work particularly well on the datasets with OOD objects; SMIYC and Synobjs. On SMIYC, our best models achieve scores above 89.0, while the next best method scores 74.9, and others score significantly lower. This indicates that the VFM pre-training is particularly helpful for out-of-distribution detection in real-world scenarios.
Method mIoU $\uparrow$ AUPR-Error $\uparrow$ AUPR-Success $\uparrow$ AUROC $\uparrow$ ECE $\downarrow$ FPR@95 $\downarrow$
DINOv2, ViT-L, 8x8 patch size, linear decoder 76.7 40.0 99.4 91.4 2.0 38.8
DINOv2, ViT-L, 16x16 patch size, linear decoder 75.9 41.2 99.5 92.3 1.7 37.8
DINOv2, ViT-g, 16x16 patch size, linear decoder 77.6 39.3 99.5 92.3 1.8 37.6
DINOv2, ViT-B, 16x16 patch size, linear decoder 71.7 43.3 99.4 92.3 2.1 40.3
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.5 45.1 99.2 91.8 2.5 44.3
PixOOD YOLO (Model Selection) 43.6 57.3 93.5 83.5 15.1 55.8
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 78.2 23.2 99.2 87.9 5.0 63.6
Model selection 73.6 44.8 99.1 91.8 9.2 40.2
PixOOD w/ ResNet-101 DeepLab 43.2 58.5 93.5 84.0 15.1 54.6
Ensemble C 73.9 47.4 99.1 92.5 52.7 34.7
Ensemble A 73.8 47.0 99.1 92.6 38.6 36.9
PixOOD w/ DeepLab Decoder 69.3 23.5 97.4 76.0 13.8 69.9
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 44.9 54.7 97.9 89.2 1.7 54.2
PixOOD 64.8 22.4 96.3 73.1 17.6 76.6
SegFormer-B5 67.4 24.1 97.2 77.2 30.7 71.9
ObsNet-R101-DLv3plus 65.3 32.1 98.5 87.8 45.6 63.4
Mask2Former-SwinB 67.2 13.2 90.4 47.0 55.1 82.8
Physically Feasible Semantic Segmentation 67.4 42.2 98.7 89.4 4.8 48.3
Table 3: Semantic metrics. Performance metrics for valid pixel predictions and their confidence, averaged across all subsets except SMIYC, computed for each method. Grayed-out methods are submitted by other participants.
Semantic metrics. Table 3 presents the performance metrics for valid pixel predictions and their confidence, averaged across all subsets except SMIYC.
We observe that the model with the highest mIoU, DINOv2 with a ViT-g/16 backbone and a Mask2Former decoder, performs relatively poorly on the other metrics. As the other metrics take into account the confidence score, this suggests that while Mask2Former is effective at predicting the correct class, it is less adept at estimating the confidence of its predictions, at least in the out-of-the-box manner in which we used it. In our setup, models with a simple linear decoder provide the best trade-off between segmentation accuracy and confidence estimation.
A similar result is observed when changing the patch size. A smaller patch size of $8\times 8$ results in better mIoU, but the other metrics are worse. This indicates that a smaller patch size allows the model to capture more fine-grained details, which improves the accuracy of the predicted class labels, but that this somehow makes the confidence scores less reliable.
Another noteworthy observation is that all our models have relatively low ECE values, indicating that they are well-calibrated, even though no explicit calibration techniques were applied. Even the DeiT-III-based model, which scores low on the overall BRAVO score, achieves a low ECE value of 1.7. Therefore, further investigation is needed to understand why the ECE values are so low.
Finally, the results suggest that more accurate models in terms of mIoU tend to be worse at identifying their own errors, as indicated by the AUPR-Error metric. However, they excel at identifying correct predictions, as shown by the AUPR-Success metric. It is possible that this happens simply because errors by accurate models are rarer, making it harder to identify them.
Overall, the results show that mIoU, which does not depend on prediction confidence, does not correlate well with the other metrics that do.
Method AUPRC $\uparrow$ AUROC $\uparrow$ FPR@95 $\downarrow$
DINOv2, ViT-L, 8x8 patch size, linear decoder 81.7 97.7 12.9
DINOv2, ViT-L, 16x16 patch size, linear decoder 76.7 97.1 15.0
DINOv2, ViT-g, 16x16 patch size, linear decoder 74.3 96.9 15.3
DINOv2, ViT-B, 16x16 patch size, linear decoder 70.6 96.6 15.1
DINOv2, ViT-S, 16x16 patch size, linear decoder 58.9 94.9 20.2
PixOOD YOLO (Model Selection) 79.0 96.5 18.2
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 84.1 98.8 4.5
Model selection 52.7 93.7 26.9
PixOOD w/ ResNet-101 DeepLab 47.1 91.7 20.8
Ensemble C 36.0 92.8 14.9
Ensemble A 32.3 91.4 17.2
PixOOD w/ DeepLab Decoder 79.0 96.5 18.2
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 30.0 86.5 38.6
PixOOD 83.0 95.2 25.7
SegFormer-B5 51.0 86.4 63.1
ObsNet-R101-DLv3plus 28.4 80.8 56.3
Mask2Former-SwinB 75.5 91.9 61.7
Physically Feasible Semantic Segmentation 26.9 83.5 74.5
Table 4: OOD metrics. Performance metrics for detecting OOD objects by identifying invalid pixels based on prediction confidence, averaged over the SMIYC and Synobjs subsets, computed for each method. Grayed-out methods are submitted by other participants.
OOD metrics. Table 4 presents the performance metrics for detecting OOD objects by identifying invalid pixels based on prediction confidence, averaged over the SMIYC and Synobjs subsets.
Surprisingly, our configuration with worst confidence estimation for valid pixels (see Table 3), DINOv2 with ViT-g/16 and Mask2Former, achieves the highest AUPRC of 84.1, the highest AUROC of 98.8, and the lowest FPR@95 of 4.5 for detecting invalid pixels. This suggests that the mask classification framework used by Mask2Former, where per-class masks are predicted separately, allows this decoder to more accurately identify which pixels belong to the mask of an in-distribution class and which do not.
Additionally, while a smaller patch size results in worse confidence estimation for valid pixels (see Table 3), it helps in identifying invalid pixels. Qualitative analyses show that the smaller patch size enables the model to better separate valid and invalid pixels, as it can capture more fine-grained details.
Finally, while scaling from ViT-L to ViT-g improves mIoU for valid pixels (see Table 3), OOD detection performance shows a noticeable degradation.
Overall, the results indicate that the models best at identifying invalid pixels are not necessarily the same ones that excel at correctly classifying valid pixels or accurately estimating their confidence for valid pixels.
5 Conclusion
This report presents the winning solution for Track 1 of the 2024 BRAVO Challenge. Our approach is based on fine-tuning Vision Foundation Models (VFMs) for semantic segmentation, leveraging the strong representations learned during pre-training. The results demonstrate that, just through better pre-training, a fine-tuned VFM is more robust under distribution shifts than complex specialist models, and can both accurately classify pixels and reliably estimate its confidence in these predictions. We evaluated our approach in several configurations, which has lead to several new observations. While we identify potential causes of these observations, future work is necessary to explore this further. Particularly, we believe it would be interesting to investigate the reason for the difference in semantic and OOD metrics with the Mask2Former decoder compared to the linear decoder. Additionally, future works could explore whether some of the more specialized out-of-distribution detection or calibration methods are still effective when used in combination with VFMs.
Acknowledgements.
This work was supported by Chips Joint Undertaking (Chips JU) in EdgeAI “Edge AI Technologies for Optimised Performance Embedded Processing” project, grant agreement No 101097300.